Which of the following pairs of equations has no solution? A}k+2j=6,k-2j=4
B}k+2j=6,k+2j=4
C}k+2j=6,k+4j=12
d}k+2j=6,k-4j=12

Answer by stanbon(75887)

Which of the following pairs of equations has no solution?
In each case check the slopes and the intercepts.
========================
A}k+2j=6,k-2j=4
k = -2j+6
k = 2j+4
Slopes different so has a solution
-------------------------------------
B}k+2j=6,k+2j=4
k = -2j+6
k = -2j+4
Slopes equal so no solution
--------------------------------------
C}k+2j=6,k+4j=12
slopes different
---------------------
d}k+2j=6,k-4j=12
slopes different
========================
Cheers,
Stan H.